If the height of a triangle is half the length of its base, and the area of the triangle is 12.25 cm squared. What is the hieght?
Answer by stanbon(75887)

If the height of a triangle is half the length of its base, and the area of the triangle is 12.25 cm squared. What is the height?
-------
base = 2x
height = x
---
Area = (1/2)base*height
-----
12.25 = (1/2)(2x)x
---
25 = 3x^2
----
x = 5/sqrt(3) = (5/3)sqrt(3)
-------------
Cheers,
Stan H.
-----------